Find the envelope of the family of the curves

1) y=3px-p^3

Where p=parameter

2) y=mx-2am-am^3

Where m=parameter


1
Expert's answer
2021-07-12T12:20:57-0400

1) The parametric equations of the envelope are defined by the system of equations


"y=3px-p^3""0=3x-3p^2"

"p=\pm\sqrt{x}"

"y=3x\sqrt{x}-x\sqrt{x}=2x\sqrt{x}"

Or


"y=-3x\sqrt{x}+x\sqrt{x}=-2x\sqrt{x}"

2) The parametric equations of the envelope are defined by the system of equations


"y=mx-2am-am^3""0=x-2a-3am^2"

"m=\pm\sqrt{\dfrac{x-2a}{3a}}"

"y=x\sqrt{\dfrac{x-2a}{3a}}-2a\sqrt{\dfrac{x-2a}{3a}}-\dfrac{x-2a}{3}\sqrt{\dfrac{x-2a}{3a}}"

"=\dfrac{3x-6a-x+2a}{3}\sqrt{\dfrac{x-2a}{3a}}"

"=\dfrac{2(x-2a)^{3/2}}{3\sqrt{3a}}"

Or


"y=-x\sqrt{\dfrac{x-2a}{3a}}+2a\sqrt{\dfrac{x-2a}{3a}}+\dfrac{x-2a}{3}\sqrt{\dfrac{x-2a}{3a}}""=\dfrac{-3x+6a+x-2a}{3}\sqrt{\dfrac{x-2a}{3a}}"

"=-\dfrac{2(x-2a)^{3/2}}{3\sqrt{3a}}"